How much 20% antifreeze solution must be added to 20 liters of 50% antifreeze solution to reduce it to 40% antifreeze?
Answer by mananth(16946)

20% antifreeze solution x liters
20 liters of 50% antifreeze solution
40% antifreeze 20 + x liters
..
0.2x+0.5*20=0.4(20+x)
0.2x+10=8 +0.4x
add -0.2x
0.2x-0.2x+10=0.4x-0.2x+8
10=0.2x+8
add -8
10-8 = 0.2x +8-8
2=0.2x
divide by 0.2
10=x
10 liters